# Quantum Oppenheimer-Snyder models in loop quantum cosmology with Lorentz term

### v1

@@ I Introduction
Recent advancements in LQG have yielded significant breakthroughs, particularly in resolving the Big Bang singularity and elucidating Hawking-Bekenstein black hole entropy [5, 6, 7]. The application of LQG principles to homogeneous, isotropic universes has established Loop Quantum Cosmology (LQC), whose hallmark achievement is replacing the classical Big Bang singularity with a quantum bounce mechanism [8, 9, 10, 11, 12, 13, 14, 15].
In classical gravitational theory, the Oppenheimer-Snyder (OS) model [16] provides a foundational framework for studying gravitational collapse through pressureless, homogeneous dust dynamics governed by Friedmann equations. While valuable, this model inherits GR’s singularity problem. Recent work has addressed this limitation by applying LQC corrections to the OS framework, resulting in the quantum Oppenheimer-Snyder (qOS) model [17], which has subsequently undergone extensive theoretical investigation [18, 19, 20].
@@
Emerging research indicates that this conventional treatment fails to fully capture LQG’s complete feature set. Crucially, full theory quantization demonstrates distinct behaviors between Euclidean and Lorentzian terms [21, 22, 23, 24]. Through Thiemann’s regularization technique, a refined LQC model incorporating the Lorentzian term has been developed. This advancement introduces significant modifications to the Friedmann equation [25, 26] manifesting an emergent de-Sitter epoch and asymmetric bounce dynamics. This raises a pivotal question: Would substituting the APS model with this Lorentzian-enhanced LQC preserve the qOS framework’s core conclusions? Our investigation addresses this through systematic derivation and analysis the physical properties of the modified qOS-LQC model.
One important physical property is that the black hole perturbation dynamics: When subjected to external perturbations, black holes exhibit a distinctive gravitational wave emission pattern that unfolds in three sequential stages. The initial burst phase, whose intensity and duration depend on the perturbation conditions, is followed by the quasinormal mode (QNM) ringdown - a hallmark feature where spacetime oscillations decay with characteristic complex frequencies. These QNM signatures, recently confirmed through groundbreaking gravitational wave detections [27, 28] have become powerful tools for probing the fundamental geometry of black holes. The ringdown profile’s precision is governed by both the fundamental mode ($n=0$) and its higher-order overtones ($n>0$) [29, 30, 31] with each overtone contributing unique information about the black hole’s structure. The final phase, known as late-time tail evolution, displays either power-law or exponential decay patterns that encode crucial information about the black hole’s properties [32], particularly regarding the nature of its Cauchy horizon [33, 34]. Consequently, the QNMs and the late-time tail, are applied to test the stability of black holes and encode important information about the black holes. To this end, we employ the finite element method and WKB approximation to investigate the QNMs, overtone, and late-time tails of scalar perturbations in quantum black holes.
@@
Concurrently, cosmological observations of universal acceleration necessitate dark energy models, where the cosmological constant $\Lambda$ remains the predominant explanation [39, 40]. In particular, the recent cosmological observations of the Universe are consistent with a de Sitter (dS) spacetime, characterized by a positive cosmological constant ($\Lambda>0$). Meanwhile, the success of the AdS/CFT correspondence has highlighted the theoretical importance of anti-de Sitter (AdS) spacetime, characterized by a negative cosmological constant ($\Lambda<0$). The AdS black holes exhibit unique thermodynamic behaviors, such as phase transitions and critical phenomena [41, 42, 43], which are not observed in asymptotically flat spacetimes. One of the most interesting black hole phase transitions is the Hawking-Page phase transition [44]. So, in this work, we focus on investigating the thermodynamic properties of black holes in both de Sitter and anti-de Sitter spacetimes.
@@ II The qOS model in LQC with Lorentz term
Here $\Delta=2\pi\sqrt{3}\gamma G\hbar$, which is the smallest non-vanishing area eigenvalue from the full theory [1, 2]. The $\gamma$ is the Barbero-Immirzi parameter. Using the black hole thermodynamics in LQG, so we could set the value being 0.2375. ${\rho_{c}}=\frac{3}{32\pi G\Delta\gamma^{2}(1+\gamma^{2})}$ is the critical energy density of this model. In the classical limit $\frac{\rho}{\rho_{c}}\ll 1$, the (2) reduce to the standard Friedmann equation. When $\rho=\rho_{c}$, the quantum bounce occurs and divides the evolution of the universe into asymmetric periods.
We should note that the modified Friedmann equation has two branches. When $f(\rho)=\frac{1+\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})}$, the corresponding modified Friedmann equation describes the pre-bounce epoch. Although the classical model does not contain a cosmological constant, the quantum geometric effects are equivalent to produce a positive effective cosmological constant of Planckian order, resulting in a de Sitter epoch. Another branch is given by $f(\rho)=\frac{1-\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})}$, which corresponds to the post-bounce epoch. This epoch is characterized by a classical universe without a cosmological constant, corresponding to the universe where we live. So in this work, we mainly focus on the post-bounce sector.
@@
To determine the unknown functions $F(r)$ and $G(r)$, we need to consider the Darmois-Israel junction condition [45, 46]. Note that the coordinates $(\tau,\tilde{r},\theta,\phi)$ describe the interior of a dust ball with a radius $0<\tilde{r}<\tilde{r}_{0}$, while the exterior is described by the coordinates $(t,r,\theta,\phi)$, with $r$ being the radial coordinate in the exterior region. To construct the complete spacetime, we need to glue the interior region described by the dust ball to the innermost region of a spherically symmetric and static exterior spacetime. In the interior region, the geodesic of a freely falling particle can be described by $\tau\mapsto(\tau,\tilde{r}_{0},\theta,\phi)$, while in the exterior region, it is described by $\tau\mapsto(t(\tau),r(\tau),\theta,\phi)$, so the single surfaces are glued by the identification $(\tau,\tilde{r}_{0},\theta,\phi)\sim(t,r,\theta,\phi)$. From the geodesic, we can derive:
$-\bigg{(}1-F\big{(}r(\tau)\big{)}\bigg{)}\dot{t}(\tau)^{2}+\bigg{(}1-G\big{(}r(\tau)\big{)}\bigg{)}^{-1}\dot{r}(\tau)^{2}=-1,\qquad-\bigg{(}1-F\big{(}r(\tau)\big{)}\bigg{)}\dot{t}(\tau)=-E.$ (4)
According to the junction condition, on the gluing surface, the induced metric and the extrinsic curvature obtained from $ds_{in}^{2}$ and $ds_{out}^{2}$ respectively are equal. This leads to:
$a(\tau)^{2}\tilde{r}_{o}^{2}=r(\tau)^{2},\qquad E^{2}\frac{1-G(r(\tau))}{1-F(r(\tau))}=1.$ (5)
Combite the (5) with the (4), we can get:
$\dot{t}(\tau)=\frac{1}{E^{2}\bigg{(}1-G\big{(}r(\tau)\big{)}\bigg{)}},\qquad\dot{a}(\tau)^{2}\tilde{r}(\tau)^{2}=\dot{r}(\tau)^{2}=G\big{(}r(\tau)\big{)}.$ (6)
Moreover, the continuity of the extrinsic curvature implies $E=1$. By substituting the modified Friedmann equation (2) into the right-hand side of Eq.(6), along the same line to that in [17], a direct calculation lead us to unambiguously fix $F(r)$ and $G(r)$. The complete form of the metric (3) is now given by:
@@
where
$F(r)=\Bigg{(}1-\frac{r^{2}f(\rho)}{\gamma^{2}\Delta}\Big{(}1-f(\rho)\Big{)}(1-\frac{\rho}{\rho_{c}})\Bigg{)},\qquad f(\rho)=\frac{1-\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})},\qquad\rho=\frac{M}{\frac{4}{3}\pi r^{3}}.$ (8)
We should note that the metric (7) is valid for
@@ III.1 Massless scalar and the effective potential
In order to map the radial region to (,+)($\infty$,$\infty$)( - ∞ , + ∞ ), we define the tortoise radial coordinate $r_{*}$ by
@@
Physically, in the event horizon, nothing can escape, so it only have incoming waves. And the metric Eq.(7) is asymptotically flat, there are no incoming waves from infinity. Thus the boundary conditions can be chosen as
@@ III.4 The QNMs results and the late-time tail
In Fig. 5 and Fig. 5, we can see that when the mass of black hole $M$ increases, the peak of the effective potential decreases. The time-domain profiles correspond to this effective potentials shown in Figure 7. In the figure, we observe that as the mass of the LQG black hole increases, the exponential decay rate decreases and the late-time tail appears later. Furthermore, the late-time tails corresponding to different black hole masses all follow an inverse power-law behavior, indicating that the black hole mass has little effect on the late-time tail.
Fig. 9 presents the effective potential $V(r_{*})$ for the Schwarzschild black hole, the black hole in the qOS model, and the qOS model with Lorentz term respectively, all evaluated with the same mass parameter $M$ and angular momentum $l$. Near the event horizon, the effective potential exhibits slight deformations, indicating modifications in the spacetime structure of the event horizon. To further explore these structural differences, we analyze the time-domain profiles in Fig. 9. As indicated in the Figure, the profile of the black hole in the previous qOS model closely matches that of the Schwarzschild black hole, whereas the exponential decay rate in the qOS model with the Lorentz term shows a noticeable decrease. The profiles of both the Schwarzschild black hole and the black hole in the qOS model with Lorentz term display a late-time inverse power-law tail, indicating that the quantum effects of the qOS model with Lorentz term have little impact on the late-time tail behavior.

### v2

@@ I Introduction
Recent advancements in LQG have yielded significant breakthroughs, particularly in resolving the Big Bang singularity and elucidating Hawking-Bekenstein black hole entropy [5, 6, 7]. The application of LQG principles to homogeneous, isotropic universes has established Loop Quantum Cosmology (LQC), whose hallmark achievement is replacing the classical Big Bang singularity with a quantum bounce mechanism [8, 9, 10, 11, 12, 13, 14, 15, 16].
In classical gravitational theory, the Oppenheimer-Snyder (OS) model [17] provides a foundational framework for studying gravitational collapse through pressureless, homogeneous dust dynamics governed by Friedmann equations. While valuable, this model inherits GR’s singularity problem. Recent work has addressed this limitation by applying LQC corrections to the OS framework, resulting in the quantum Oppenheimer-Snyder (qOS) model [18], which has subsequently undergone extensive theoretical investigation [19, 20, 21, 22].
@@
Emerging research indicates that this conventional treatment fails to fully capture LQG’s complete feature set. Crucially, full theory quantization demonstrates distinct behaviors between Euclidean and Lorentzian terms [23, 24, 25, 26]. Through Thiemann’s regularization technique, a refined LQC model incorporating the Lorentzian term has been developed. This advancement introduces significant modifications to the Friedmann equation [27, 28, 29] manifesting an emergent de-Sitter epoch and asymmetric bounce dynamics. This raises a pivotal question: Would substituting the APS model with this Lorentzian-enhanced LQC preserve the qOS framework’s core conclusions? Our investigation addresses this through systematic derivation and analysis the physical properties of the modified qOS-LQC model.
One important physical property is that the black hole perturbation dynamics: When subjected to external perturbations, black holes exhibit a distinctive gravitational wave emission pattern that unfolds in three sequential stages. The initial burst phase, whose intensity and duration depend on the perturbation conditions, is followed by the quasinormal mode (QNM) ringdown - a hallmark feature where spacetime oscillations decay with characteristic complex frequencies [30]. These QNM signatures, recently confirmed through groundbreaking gravitational wave detections [31, 32] have become powerful tools for probing the fundamental geometry of black holes. The ringdown profile’s precision is governed by both the fundamental mode ($n=0$) and its higher-order overtones ($n>0$) [33, 34, 35] with each overtone contributing unique information about the black hole’s structure. The final phase, known as late-time tail evolution, displays either power-law or exponential decay patterns that encode crucial information about the black hole’s properties [36], particularly regarding the nature of its Cauchy horizon [37, 38]. Consequently, the QNMs and the late-time tail, are applied to test the stability of black holes and encode important information about the black holes. To this end, we employ the finite element method and WKB approximation to investigate the QNMs, overtone, and late-time tails of scalar perturbations in quantum black holes.
@@
Concurrently, cosmological observations of universal acceleration necessitate dark energy models, where the cosmological constant $\Lambda$ remains the predominant explanation [43, 44]. In particular, the recent cosmological observations of the Universe are consistent with a de Sitter (dS) spacetime, characterized by a positive cosmological constant ($\Lambda>0$). In this context, the physical and thermodynamic properties of de Sitter black holes have also been extensively studied [45, 46]. Meanwhile, the success of the AdS/CFT correspondence has highlighted the theoretical importance of anti-de Sitter (AdS) spacetime, characterized by a negative cosmological constant ($\Lambda<0$). The AdS black holes exhibit unique thermodynamic behaviors, such as phase transitions and critical phenomena [47, 48, 49], which are not observed in asymptotically flat spacetimes. One of the most interesting black hole phase transitions is the Hawking-Page phase transition [50]. In this work, we focus on investigating the thermodynamic properties in both de Sitter and anti-de Sitter spacetimes.
@@ II The qOS model in LQC with Lorentz term
Here $\Delta=2\pi\sqrt{3}\gamma G\hbar$, which is the smallest non-vanishing area eigenvalue from the full theory [1, 2]. The $\gamma$ is the Barbero-Immirzi parameter. Using the black hole thermodynamics in LQG, so we could set the value being 0.2375. ${\rho_{c}}=\frac{3}{32\pi G\Delta\gamma^{2}(1+\gamma^{2})}$ is the critical energy density of this model.
We should note that the modified Friedmann equation possesses two distinct branches. The first branch is $f(\rho)_{+}=\frac{1+\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})}$, though it has no cosmological constant added at the classical evolution. However, its quantum geometric effects are equivalent to a positive effective cosmological constant of Planckian order [23]. The effective cosmological constant is $\Lambda_{eff}=\frac{3}{(1+\gamma^{2})^{2}\Delta}$, resulting in an de-Sitter epoch. When $\rho\ll\rho_{c}$, the $f(\rho)_{+}$ branch of the modified Friedmann equation reduces to
$H^{2}_{+}=\frac{8\pi G_{\alpha}}{3}\rho+\frac{\Lambda_{eff}}{3}.$ (3)
where $G_{\alpha}=\frac{1-5\gamma^{2}}{\gamma^{2}+1}G$. This differs from the classical case, as it includes both a positive effective cosmological constant and a modified Newton’s constant.
The second branch is given by $f(\rho)_{-}=\frac{1-\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})}$. When $\rho\ll\rho_{c}$, it reduces to
$H^{2}_{-}=\frac{8\pi G}{3}\rho.$ (4)
This branch corresponds to a classical universe in which no effective cosmological constant appears.
@@
It is worthy noting that in the cosmological case, the $f(\rho)_{+}$ branch is not suitable for describing our universe, and therefore the bounce proceeds from the $f(\rho)_{+}$ branch to the $f(\rho)_{-}$ branch. Conversely, in the collapse case, the $f(\rho)_{-}$ branch provides a more appropriate description of stellar collapse, so the bounce proceeds from the $f(\rho)_{-}$ branch to the $f(\rho)_{+}$ branch also sounds. Since in both scenarios we are primarily interested in post-bounce branch, the resulting exterior solutions differ accordingly. A detailed explanation will be provided below.
II.1 The vacuum exterior for the $f(\rho)_{-}$ branch
As Eq. (3) shows, the emergence of the effective cosmological constant and the modified Newton’s constant is not consistent with current cosmological observations. Therefore, it is not suitable for describing our expanding universe. In the cosmological scenario, the bounce transitions from the $f(\rho)_{+}$ branch to the $f(\rho)_{-}$ branch, where the universe corresponding to the $f(\rho)_{-}$ branch is the one we inhabit. Accordingly, we focus on the $f(\rho)_{-}$ in the following analysis.
Having fixed the branch of the modified Friedmann equation, we can now determine the metric outside the dust ball region. To determine the unknown functions $F(r)$ and $G(r)$, we need to consider the Darmois-Israel junction condition [51, 52]. We first consider the radial coordinate of the surface of the dust ball, denoted by $\tilde{r_{0}}$. Then, the hypersurface induced from Eq.(1) is
$ds_{in}^{2}|_{\Sigma}=-d\tau^{2}+a(\tau)^{2}\tilde{r_{0}}^{2}d\Omega^{2}.$ (6)
We assume that every point on the surface of the dust ball shares the same proper time $\tau$, which evolves as the collapse proceeds. For convenience, we use $\tau$ to parameterize the hypersurface in the vacuum region outside the dust ball. Then, we have
$ds_{out}^{2}|_{\Sigma}=-\bigg(F\big(r(\tau)\big)\dot{t}(\tau)^{2}-G\big(r(\tau)\big)^{-1}\dot{r}(\tau)^{2}\bigg)d\tau^{2}+r(\tau)^{2}d\Omega^{2}.$ (7)
To construct the complete spacetime, we match the interior region of the dust ball to the exterior region of a spherically symmetric, static spacetime using the Darmois–Israel junction conditions. The first junction condition requires the metric to be continuous across the boundary. Consequently, at the boundary, the metric components satisfy
$1=\bigg(F\big(r(\tau)\big)\dot{t}(\tau)^{2}-G\big(r(\tau)\big)^{-1}\dot{r}(\tau)^{2}\bigg),\quad\quad a(\tau)=r(\tau).$ (8)
According to the second Darmois–Israel junction condition, the extrinsic curvature must be continuous across the boundary. In the spacetime of the interior dust ball, the non-vanishing components are
$K^{in}_{\theta\theta}=a(\tau)\tilde{r_{0}},\quad K^{in}_{\phi\phi}=a(\tau)\tilde{r_{0}}\sin^{2}\theta.$ (9)
For the exterior vacuum spacetime, the non-zero components of the extrinsic curvature are
$K^{out}_{\theta\theta}=rF(r(\tau))\dot{t}\sqrt{F(r(\tau))^{-1}G(r(\tau))},\quad K^{out}_{\phi\phi}=rF(r(\tau))\dot{t}\sqrt{F(r(\tau))^{-1}G(r(\tau))}\sin^{2}\theta.$ (10)
Since the exterior spacetime is assumed to be static, $\xi^{a}=(\frac{\partial}{\partial t})^{a}$ is a Killing vector, which allows us to define the conserved energy
$\xi_{\mu}u^{\mu}=g^{out}_{\mu\nu}\xi^{\nu}u^{\mu}=F\big(r(\tau)\big)\dot{t}(\tau)=-E,$ (11)
where $u^{\mu}=(\dot{t},\dot{r},0,0)$.
Combining Eq. (9) and Eq. (10), and using Eq. (11), we obtain
$a(\tau)\tilde{r_{0}}=rF(r(\tau))\dot{t}(\tau)\sqrt{F(r(\tau))^{-1}G(r(\tau))}=rE\sqrt{F(r(\tau))^{-1}G(r(\tau))}.$ (12)
Using Eq. (12) together with the metric continuity condition Eq. (8), we find
$F(r(\tau))=E^{2}G(r(\tau)),\quad G(r(\tau))=1-\dot{r}^{2}.$ (13)
Combining Eq. (13) with the modified Friedmann equation Eq. (2), we obtain
$G(r(\tau))=1-H^{2}r^{2}=1-\frac{r^{2}}{\gamma^{2}\Delta}f(\rho)\big(1-f(\rho)\big)\big(1-\frac{\rho}{\rho_{c}}\big).$ (14)
For simplicity and without loss generality, we can set $E=1$, which gives $F(r)=G(r)$. The exterior metric then takes the form
@@ II.1 The vacuum exterior for the $f(\rho)_{-}$ branch
$F(r)$ $=1-\frac{r^{2}}{2\gamma^{2}(1+\gamma^{2})\Delta}\left(1-\dfrac{1-\sqrt{1-\dfrac{3M}{4\pi r^{3}\rho_{c}}}}{2(1+\gamma^{2})}\right)\left(1-\sqrt{1-\dfrac{3M}{4\pi r^{3}\rho_{c}}}\right)\left(1-\dfrac{3M}{4\pi r^{3}\rho_{c}}\right).$ (15)
It should be noted that the metric (15) is asymptotically flat, and is valid for
@@
II.2 The vacuum exterior for the $f(\rho)_{+}$ branch
In the classical Oppenheimer–Snyder (OS) model, a homogeneous dust ball collapses inevitably to a singularity. In contrast, in the de Sitter OS model [53] contracts to a minimal radius and undergoes a bounce, followed by expansion. Studying the physical properties and formation thresholds of primordial black holes formed from gravitational collapse within the LQC framework provides a valuable probe of the quantum nature of gravity and its potential connection to dark matter [54]. Therefore, considering gravitational collapse in this context is both important and meaningful.
In the context of stellar collapse, the interior of the dust ball is described by the metric
$ds_{-}^{2}=-d\tau^{2}+a(\tau)^{2}(d\tilde{r}^{2}+\tilde{r}^{2}d\Omega^{2}),$ (17)
The dynamics of $a(\tau)$ are governed by the $f(\rho)_{-}$ branch of the modified Friedmann equation, given by
$H_{-}^{2}=(\frac{\dot{a}}{a})^{2}=\frac{1}{2\gamma^{2}(1+\gamma^{2})\Delta}\left(1-\dfrac{1-\sqrt{1-\dfrac{3M}{4\pi r^{3}\rho_{c}}}}{2(1+\gamma^{2})}\right)\left(1-\sqrt{1-\dfrac{3M}{4\pi r^{3}\rho_{c}}}\right)\left(1-\dfrac{3M}{4\pi r^{3}\rho_{c}}\right).$ (18)
The dust ball starts from a nearly static configuration at large $r$, contracts gradually, and crosses the past event horizon. It continues collapsing until it reaches the minimal radius $r_{b}$ where the bounce occurs. After the bounce, the evolution transitions to the $f(\rho)_{+}$ branch, representing the expanding phase, described by the metric
$ds_{+}^{2}=-d\tau^{2}+a(\tau)^{2}(d\tilde{r}^{2}+\tilde{r}^{2}d\Omega^{2}),$ (19)
with the scale factor $a(\tau)$ governed by the $f(\rho)_{+}$ branch:
$H_{+}^{2}=(\frac{\dot{a}}{a})^{2}=\frac{1}{2\gamma^{2}(1+\gamma^{2})\Delta}\left(1-\dfrac{1+\sqrt{1-\dfrac{3M}{4\pi r^{3}\rho_{c}}}}{2(1+\gamma^{2})}\right)\left(1+\sqrt{1-\dfrac{3M}{4\pi r^{3}\rho_{c}}}\right)\left(1-\dfrac{3M}{4\pi r^{3}\rho_{c}}\right).$ (20)
The dust ball then expands outward, crosses the future cosmological horizon, and continues to infinity.
It should be noted that only the $f(\rho)_{-}$ branch reproduces the classical Oppenheimer–Snyder collapse in the $\rho\ll\rho_{c}$, while the $f(\rho)_{+}$ branch does not, as discussed previously. Consequently, the big bounce corresponds to a transition from the $f(\rho)_{-}$ branch to the $f(\rho)_{+}$ branch, and we are primarily interested in post-bounce branch.
Once the appropriate branch of the modified Friedmann equation is identified, the exterior metric of the dust ball can be determined following the same procedure as outlined before, resulting in the metric
$ds_{out}^{2}$ $=-F(r)dt^{2}+F(r)^{-1}dr^{2}+d\Omega^{2},$
$F(r)$ $=1-\frac{r^{2}}{2\gamma^{2}(1+\gamma^{2})\Delta}\left(1-\frac{1+\sqrt{1-\frac{3M}{4\pi r^{3}\rho_{c}}}}{2(1+\gamma^{2})}\right)\left(1+\sqrt{1-\frac{3M}{4\pi r^{3}\rho_{c}}}\right)\left(1-\frac{3M}{4\pi r^{3}\rho_{c}}\right).$ (21)
@@ III QUASINORMAL MODES
As mentioned earlier, the vacuum exterior of the $f(\rho)_{-}$ branch corresponds to a spacetime containing a black hole, whereas the $f(\rho)_{+}$ branch admits only a cosmological horizon. Moreover, the vacuum exterior of the $f(\rho)_{-}$ branch appears to be more consistent with cosmological observations. Therefore, in this section, we restrict our analysis to the $f(\rho)_{-}$ branch.
@@ III.1 Massless scalar and the effective potential
In order to map the radial region to $(-\infty,+\infty)$, we define the tortoise radial coordinate $r_{*}$ by
@@
Physically, in the event horizon, nothing can escape, so it only have incoming waves. And the metric Eq.(LABEL:eq:_new_metric) is asymptotically flat, there are no incoming waves from infinity. Thus the boundary conditions can be chosen as
@@ III.4 The QNMs results and the late-time tail
In Fig. 5 and Fig. 5, we can see that when the mass of black hole $M$ increases, the peak of the effective potential decreases. The time-domain profiles correspond to this effective potentials shown in Figure 7. In the figure, we observe that as the mass of the LQG-inspired black hole increases, the exponential decay rate decreases and the late-time tail appears later. Furthermore, the late-time tails corresponding to different black hole masses all follow an inverse power-law behavior, indicating that the black hole mass has little effect on the late-time tail.
Fig. 9 presents the effective potential $V(r_{*})$ for the Schwarzschild black hole, the black hole in the qOS model, and the qOS model with Lorentz term respectively, all evaluated with the same mass parameter $M$ and angular momentum $l$. Near the event horizon, the effective potential exhibits slight deformations, indicating modifications in the spacetime structure of the event horizon. To further explore these structural differences, we analyze the time-domain profiles in Fig. 9. As indicated in the Figure, the profile of the black hole in the previous qOS model closely matches that of the Schwarzschild black hole, whereas the exponential decay rate in the qOS model with the Lorentz term shows a noticeable decrease. It should be noted that the deformation induced by quantum effects is minimal compared to the classical case. The profiles of both the Schwarzschild black hole and the black hole in the qOS model with Lorentz term display a late-time inverse power-law tail, indicating that the quantum effects of the qOS model with Lorentz term have little impact on the late-time tail behavior.